# Safe and Generalized end-to-end Autonomous Driving System with Reinforcement Learning and Demonstrations

### v3

@@ 3.2 Policy Learning
In the second stage of the model, we employ $\mathcal{M}_{\text{expert}}$ to initialize a model trained by deep reinforcement learning policies, which greatly reduces the sampling complexity of the deep RL policy intelligent vehicle. The training loss of the actor model in the second stage is as follows:
${\rm\mathcal{M}_{actor}(\phi,\eta)}=\mathbb{E}\Big{(}\sum_{i=1}^{N}\sum_{\tau=t}^{t+H}\left(V_{i,\tau}^{\lambda}\right)\Big{)}+k\ln\pi^{*}_{\eta}(a_{t}^{\prime}|s_{t}^{\prime})$ (5)
Where $k$ represents the balance between the cloning policy and deep reinforcement learning policy, and is set as a constant based on empirical data. Then we can train the critic to regress the TD($\lambda$) (Sutton & Barto, 2018) target return via a mean squared error loss:
$\mathcal{M}_{\text{critic }}(\xi)=\mathbb{E}\Big{[}\sum_{i=1}^{N}\sum_{\tau=t}^{t+H}\frac{1}{2}\left(v_{\xi}\left(s_{i,\tau}\right)-V_{i,\tau}^{\lambda}\right)^{2}\Big{]}$ (6)
The $v_{\xi}\left(s_{\tau}\right)$ is designed to forecast the cumulative discounted sum of future rewards achievable by the actor, starting from an initial state $s_{\tau}$. $V_{\tau}^{\lambda}$ can be defined as follow:
$V_{\tau}^{\lambda}=(1-\tau)v_{\xi}(s_{\tau+1})+\lambda V_{\tau+1}^{\lambda},\quad\tau\textless{t+H}$ (7)
Where $\xi$ denote the parameters of the critic network and $H$ is the prediction horizon. Finally, the loss function for the second stage training of the whole model is as follows:
$\underset{\psi,\xi,\phi,\theta,\omega,\eta}{\rm min}\alpha_{0}{\rm\mathcal{M}_{critic}(\xi)}-\alpha_{1}{\rm\mathcal{M}_{actor}(\phi,\eta)}-\alpha_{2}{\rm\mathcal{M}_{model}(\theta,\psi,\omega)}$ (8)
We jointly optimize the parameters of model loss $\psi,\theta,\omega$, critic loss $\xi$ and actor loss $\phi,\eta$, where $\alpha_{0},\alpha_{1},\alpha_{2}$ are coefficients for different components.
3.3 Safety constraint
Figure 4: Randomly sample sensor inputs Lidar_noground $o_{1},o_{2},\cdots,o_{15}$, and then our model can imagine driving behaviors $\hat{o}_{16},\hat{o}_{17},\cdots,\hat{o}_{30}$. The results show that our model can make precise and diverse predictions with no mode mixup and less blur.
To keep the intelligent vehicle running safely and smoothly in complex environments, we add safety and robustness constraints to the reward function of the algorithm(Chen et al., 2019b), including time to collision and smooth steering.
(1)Front time to collision. When around vehicles are within the distance of ego vehicle (our agent vehicle) head in our setting, then we can calculate the front time to collision between ego vehicle and around vehicles. Firstly, the speed and steering vector $(s_{\tau},a_{\tau})\in\mathbb{S}$ of the ego vehicle are defined, where $s_{\tau}$ represents the angle vector of vehicle steering and $a_{\tau}$ represents the acceleration vector of the vehicle in local coordinate system. Secondly, two waypoints closest to the current ego vehicle are selected from the given navigation routing as direction vectors $\vec{w_{p}}$ for the entire route progression, where $\rightarrow$ indicates a vector in world coordinates. The position vectors for both ego vehicle and around vehicles are represented by $(\vec{x_{t}}^{*},\vec{y_{t}}^{*})$, respectively. Finally, $\delta_{e}$ and $\delta_{a}$ representing angles between position vectors for ego vehicle and around vehicles with respect to $\vec{w_{p}}$ are calculated respectively.
$\vec{w_{p}}=\big{[}(\frac{\vec{w_{t+1}^{x}}-\vec{w_{t}^{x}}}{2})-(\vec{w_{t}^{x}}),(\frac{\vec{w_{t+1}^{y}}-\vec{w_{t}^{y}}}{2})-(\vec{w_{t}^{y}})\big{]}$ (9)
$\delta_{e}=\frac{[\vec{v_{t}^{x*}},\vec{v_{t}^{y*}}]\cdot\vec{w_{p}}}{\|\vec{v_{t}^{x*}},\vec{v_{t}^{y*}}\|_{2}\quad\|\vec{w_{p}}\|_{2}},\delta_{a}=\frac{[\vec{v_{t}^{x}},\vec{v_{t}^{y}}]\cdot\vec{w_{p}}}{\|\vec{v_{t}^{x}},\vec{v_{t}^{y}}\|_{2}\quad\|\vec{w_{p}}\|_{2}}$
Where, $l$ is the length of the set of waypoints $\mathbb{W}$ stored. The variable $t\in\tau$, and $\vec{w_{t}^{x}}\in\mathbb{W}_{1}$ represents the $x$ coordinate of the first navigation point closest to the intelligent vehicle on its current route at time $t$. Similarly, $\vec{w_{t+1}^{x}}\in\mathbb{W}_{2}$. Furthermore, it is possible to calculate the $F_{ttc}$ as follows:
$F_{ttc}=\frac{\|\vec{x_{t}}-\vec{x_{t}^{*}},\vec{y_{t}}-\vec{y_{t}^{*}}\|_{2}}{\left|\|\vec{v_{t}^{x*}},\vec{v_{t}^{y*}}\|_{2}sin(\delta_{e})-\|\vec{v_{t}^{x}},\vec{v_{t}^{y}}\|_{2}sin(\delta_{a})\right|}$ (10)
$&\text{min}(z_{\tau},c_{\tau}+1.0),\\ &\nu_{g}\leq(c_{lttc}-1.5)\text{ and }\mu_{a}\leq(c_{\tau}-0.5).\\ &\text{min}(z_{\tau},c_{\tau}-1.8),\\ &\nu_{g}\leq(c_{\tau}-3.0)\text{ and }\mu_{a}\leq(c_{\tau}-2.0).\\ &\text{min}(z_{\tau},c_{\tau}-3.0),\\ &\nu_{g}\leq(c_{\tau}-3.5)\text{ and }\mu_{a}\leq(c_{\tau}-3.0).\\ \end{aligned}\right.\vspace{-0.2cm}$ (11)
@@ 4 Experiment
Figure 5: The road networks of the CARLA include routes for Town01, Town02, Town03, Town04, Town05, and Town06, as well as more complex scenarios in Town03, Town04, and Town05
The model training in this paper is deployed on systems including Python=3.6, Pytorch-gpu=1.5.0, GeForce RTX-3090 and CARLA=0.9.10 running on Linux. In order to evaluate the performance of our SGADS in addressing challenges related to sampling efficiency, safety, generalization and prediction for future driving trajectories, we respectively designed corresponding experiments to answer each of these questions. Additionally, we formulated experiments involving 12 different sets of input data to evaluate the influence on the safety of autonomous vehicles.
Table 1: In the training stage, the steps and times for all methods are compared under different baseline average collision distances in Town03(max episodes length=500), the number of vehicles is 200, and calculating Average Driving Distance(Avg Dis) with 10 episodes. Particularly, +$\infty$+ ∞ means the situation where the training time has exceeded 250hours and the model has still not reached the baseline.
Method Input Avg dis=50m Avg dis=150m Avg dis=200m
$\downarrow$ $\downarrow$ $\downarrow$ times $\downarrow$ $\downarrow$ times $\downarrow$
DDPG Lidar +$\infty$+ ∞ +$\infty$+ ∞ +$\infty$+ ∞ +$\infty$+ ∞ +$\infty$+ ∞ +$\infty$+ ∞
SAC Lidar +$\infty$+ ∞ +$\infty$+ ∞ +$\infty$+ ∞ +$\infty$+ ∞ +$\infty$+ ∞ +$\infty$+ ∞
TD3 Lidar $\geq$161 $\geq$192h +$\infty$+ ∞ +$\infty$+ ∞ +$\infty$+ ∞ +$\infty$+ ∞
DQN Lidar $\geq$163 $\geq$53h +$\infty$+ ∞ +$\infty$+ ∞ +$\infty$+ ∞ +$\infty$+ ∞
Latent_SAC Lidar $\geq$167 $\geq$40h $\geq$300 $\geq$212h +$\infty$+ ∞ +$\infty$+ ∞
SGADS Lidar $\geq$21 $\geq$1.3h $\geq$121 $\geq$20h $\geq$321 $\geq$48h
SGADS Lidar_ng $\geq$15 $\geq$0.9h $\geq$139 $\geq$9h $\geq$211 $\geq$38h
4.1 Training and Evaluation Setting
4.1.1 Traning environment

### v8

@@ 3.6 Safety constraint
3.7 Augmenting RL policy with demonstrations
Though, NFRL can significantly reduce complexity, and reward design based on safety constraints can enhance safety. Demonstrations can mitigate the need for painstaking reward shaping, guide exploration, further reduce sample complexity, and help generate robust, natural behaviors. We propose the demonstration augmented RL agent method which incorporates demonstrations into NFRL agent in two ways:
(1) Pretraining with behavior cloning. We use behavior cloning to provide a policy $\pi^{*}$ via expert demonstrations and then to train a model $\mathcal{M}_{\text{expert}}$ with some expert ability.
$\mathcal{M}_{\text{expert}}={\underset{\xi}{\rm maximize}}\sum_{(s^{\prime},a^{\prime})\in\pi^{*}(\mathcal{D}_{e})}\ln\pi^{*}_{\xi}(a_{\tau}^{\prime}|s_{\tau}^{\prime})$ (12)
where $\mathcal{D}_{e}$ is a human expert dataset obtained from driving G29 steering. For detailed information on all of the expert dataset, please refer to Appendix  A.2.
(2) RL fine-tuning with augmented loss: we employ $\mathcal{M}_{\text{expert}}$ to initialize a model trained by deep RL policies, which reduces the sampling complexity of the deep RL policy. The training loss of the actor model as follows:
${\rm\mathcal{M}_{\hat{\text{actor}}}(\phi,\xi)}=\mathcal{M}_{\text{actor}}(\phi)+k\ln\pi^{*}_{\xi}(a_{\tau}^{\prime}|s_{\tau}^{\prime}),(a_{\tau}^{\prime},s_{\tau}^{\prime})\in\mathcal{D}_{e}$ (13)
where $k$ represents the balance between the behavior cloning policy and NFRL policy, and is set as a constant based on empirical data. We only changed the actor model of NFRL, and the optimization of the other parts is exactly the same.
Figure 3: Randomly sample sensor inputs Lidar_noground $o_{1},o_{2},\cdots,o_{15}$, and then our model can imagine driving behaviors $\hat{o}_{16},\hat{o}_{17},\cdots,\hat{o}_{30}$. The results show that, compared with Dreamer using ground truth data, our NFRL model is more accurate and diverse, with no mode mixup and less blur.
@@ 4 Experiment
4.1 Experiment setup
Models were trained on an NVIDIA RTX 3090 GPU using Python 3.8. Our experiments were conducted on a benchmark called Gym-carla, a third-party environment for OpenAI gym that is used with the CARLA simulatordosovitskiy2017carla . In our experiments, the NFRL series models and baseline methods were trained in Town03 (random) and evaluated across four scenarios: Town03, Town04, Town05, and Town06. These scenarios are abbreviated as Town03-Town06, encompassing both random and roundabout modes. Town03 simulates a realistic urban environment with diverse features such as tunnels, intersections, roundabouts, curves, and turnaround bends. Descriptions of other maps are provided in Appendix  A.1. Here, "random" refers to randomly selected intersections and driving scenarios, while "roundabout" focuses specifically on roundabout intersections. We present the hyperparameter settings for the methods in Appendix  A.6. Particularly we contribute a dataset of human expert $\mathcal{D}_{expert}$ via G29 steering wheel, with further details provided in Appendix  A.2 description.
4.2 Measure Driving Performance Metrics
In the Gym-Carla benchmark, an episode terminates under any of the following conditions: the number of collisions exceeds one, the maximum number of time steps is reached, the destination is reached, the cumulative lateral deviation from the lane exceeds 10 meters, or the vehicle remains stationary for 50 seconds. EGADS is an end-to-end autonomous driving system. We implemented our trained model on an autonomous vehicle for urban navigation, assessing performance through five standard metrics: Off-road Rate (OR), Episode Completion Rate (ER), Average Safe Driving Distance (ASD), Average Reward (AR) using the reward function from Chen et.al chen2019model that accounts for driving dynamics (yaw, collisions, speeding, and lateral velocity), and Driving Score (DS) - a composite metric calculated as DS = ER × AR in accordance with CARLA Leaderboard standards. During model selection, we focused on checkpoints that simultaneously optimized DS and AR, while implementing the remaining metrics (ER, OR, AR, ASD) based on the methodology from Gao et.al gao2024enhance and Tang et.al tangvlascd ; tang2024efficient . For detailed information on all of the above metrics, please refer to Appendix  A.5.
4.3 Comparison settings
In order to evaluate the performance of our autonomous driving system more effectively, we have conducted various comparisons with existing methods such as DDPG lillicrap2015continuous , SAC haarnoja2018soft , TD3 fujimoto2018addressing , DQN mnih2015human , Latent_SAC chen2021interpretable , Dreamer hafner2019dream , CQL kumar2020conservative . We decomposed EGADS into three components: NFRL, SC (safety constraint) and augmenting RL policy with demonstrations. We decomposed EGADS into three components: NFRL, Safety Constraint (SC) and augmenting RL policy with demonstrations (Demo). We then conducted evaluations using four comparison settings, NFRL, NFRL+SC, BC+Demo, NFRL+SC+Demo. BC+Demo indicates the use of behavioral cloning to imitate the expert dataset, while NFRL+SC+Demo involves using expert datasets to augment the NFRL policy combined with SC.
Table 1: In training, all methods were compared under different RL baselines in Town03 (random), with episodes of 500 steps. $\infty$ indicates failure to reach the baseline within the maximum tested runtime of 250 GPU hours.
Method ASD=50m ASD=100m
$downarrow episodes\downarrow times \downarrow$
DDPG $\infty ++\infty$ $\infty ++\infty$
SAC $\infty ++\infty$ $\infty ++\infty$
TD3 $\geq$161 $\geq$192h $\infty ++\infty$
DQN $\geq$163 $\geq$53h $\infty ++\infty$
Latent_SAC $\geq$167 $\geq$43h $\geq$352 $\geq$105h
Dreamer $\infty ++\infty$ $\infty ++\infty$
NFRL(our) $\geq$141 $\geq$21h $\geq$121 $\geq$65h
Table 2: In training, all methods were compared under different NFRL baselines in Town03 (random), with episodes of 500 steps. $\infty$ indicates failure to reach the baseline within the maximum tested runtime of 250 GPU hours.
Method ASD=50m ASD=100m ASD=200m
[…]
NFRL $\geq$141 $\geq$21h $\geq$121 $\geq$65h $\infty ++\infty$
NFRL+SC $\geq$71 $\geq$12h $\geq$301 $\geq$40h $\geq$1100 $\geq$146h
NFRL+SC+Demo $\geq$21 $\geq$1.3h $\geq$58 $\geq$3h $\geq$321 $\geq$48h
4.4 Results on trajectory prediction
In order to accurately evaluate our model prediction of driving actions for intelligent vehicle, this problem can be viewed as a special POMDP problem with the reward value maintained at 0. As shown in Figure 3, the comparison with ground-truth data demonstrates that our NFRL model achieves higher accuracy and greater diversity than Dreamer, with no mode collapse and significantly reduced blurring effects. We provide additional results on predictions of future driving actions for NFRL in Appendix  A.8.
4.5 How to reduce sampling complexity ?
To evaluate the sampling complexity of different methods, we used the average ASD as the test threshold and set three distinct checkpoints at 50m, 100m, and 200m. We measured the GPU hours required for each method to reach the corresponding ASD threshold, with a maximum testing duration capped at 250 GPU hours, as shown in Tables 1 and 2. Notably, in Table 1, although different methods require varying numbers of episodes to reach the ASD threshold, the actual time consumed differs significantly. This is because each episode has a fixed length of 500 steps. Some methods remain stationary for most of the episode, yet the episode does not terminate early, leading to prolonged total runtime. In contrast, other methods may collide or deviate from the lane, triggering early termination of the episode.
As shown in Table 1, our proposed NFRL method significantly improves training time efficiency, achieving at least a 2-fold acceleration in reaching the 50-meter and 100-meter baselines compared to existing reinforcement learning methods. However, due to frequent collision issues observed in experiments, the method fails to surpass the 150-meter baseline. To address this limitation, we innovatively design a reward function incorporating Safety Constraints (SC). Experimental results, as presented in Table 2, show that the enhanced NFRL+SC method not only successfully achieves the 200-meter baseline but also improves training efficiency by at least 1.5 times compared to the original NFRL method. To further optimize performance, we introduce expert datasets for fine-tuning. Experimental data indicate that the NFRL+SC+Demo method achieves a remarkable 3-fold improvement in training efficiency over the NFRL+SC method when reaching the 200-meter baseline.
The performance improvements are primarily driven by three key mechanisms: (1) The NFRL framework employs Normalizing Flow technology to reconstruct training data distributions, aligning them more closely with real-world driving scenarios. This technique enables both accurate future trajectory prediction and comprehensive coverage of possible trajectories across diverse driving situations. Such high-quality data representation allows the model to rapidly learn correct behavioral patterns. (2) The Safety Constraint (SC) module dynamically limits the policy exploration scope to safe regions, thereby minimizing costly divergent behaviors. (3) Demonstration data accelerates reward function discovery by injecting domain-specific prior knowledge. This co-design enables EGADS to achieve efficient convergence in complex autonomous driving scenarios, establishing it as a paradigm for sample-efficient reinforcement learning.
Table 3: Performance Comparison Across multiple Towns (Trained in Town03, Evaluated in Town04-Town06, hereinafter referred to as T04-T06)
Method DS $\uparrow$ AR ($f_{1}$) $\uparrow$ EC (%) $\uparrow$ OR (%) $\downarrow$ ASD (m) $\uparrow$
T04 T05 T06 T04 T05 T06 T04 T05 T06 T04 T05 T06 T04 T05 T06
DDPG -0.10 -0.01 -0.08 -10.01 -10.1 -10.02 0.00 0.01 0.00 - - - 0.00 0.00 0.00
DQN 17.50 60.67 69.09 76.37 174.89 206.66 11.38 15.84 16.34 11.83 11.83 15.26 20.29 31.01 36.83
TD3 -15.89 -2.24 -25.62 -131.36 -84.30 -195.60 9.18 8.16 5.82 33.32 16.94 16.02 6.17 10.05 4.50
SAC -20.56 -14.89 -15.02 -14.08 -18.92 -16.67 4.95 69.07 85.60 0.00 0.00 0.00 6.71 6.07 8.03
L_SAC 102.61 110.66 21.52 170.79 8.70 145.77 15.09 12.78 12.21 1.05 4.96 4.64 15.97 21.24 42.15
Dreamer -0.01 -0.03 -0.03 -15.10 -15.10 -15.20 0.00 0.12 0.20 - - - 0.01 0.01 0.00
NFRL (base) 326.78 390.54 431.44 1509.90 785.92 947.26 15.81 22.61 29.59 30.88 12.05 16.50 220.18 123.24 143.61
Table 4: Evaluation results for different methods in CARLA Town03 (random) and Town03 (roundabout): we denote RND as random and RBT as roundabout. For a fair comparison, all reward functions are in the form of $f_{1}$. Particularly, - indicates that valid data could not be obtained because the episode completion rate for this method is close to 0.
Method DS $\uparrow$ AR ($f_{1}$) $\uparrow$ EC(%) $\uparrow$ OR(%) $\downarrow$ ASD(m) $\uparrow$
RND RBT RND RBT RND RBT RND RBT RND RBT
DDPG 0.11-0.11 0.08-0.08 10.01-10.01 10.02-10.02 $0.01$ $0.00$ - - $0.00$ $0.00$
DQN $30.64$ $36.33$ $86.50$ $121.24$ $17.02$ $16.42$ $8.52$ $11.83$ $21.68$ $26.27$
TD3 $2.40$ 6.60-6.60 18.15-18.15 129.52-129.52 $6.91$ $4.07$ $51.53$ $49.32$ $7.51$ $3.12$
SAC 7.57-7.57 20.56-20.56 19.90-19.90 24.74-24.74 63.76 67.95 0.00 0.65 $6.27$ $6.71$
L_SAC $125.95$ $31.59$ $161.24$ $84.13$ $11.98$ $10.50$ $14.72$ $1.14$ $31.31$ $13.87$
Dreamer 0.03-0.03 0.02-0.02 15.12-15.12 15.12-15.12 $0.01$ $0.02$ - - $0.01$ $0.01$
NFRL (base) 170.03 48.73 424.60 249.17 $24.04$ $10.88$ $20.93$ $18.11$ 72.16 46.27
4.6 Ablation study
In the ablation study of the EGADS system, we evaluated the contributions of each module in cross-domain scenarios (evaluated in Town04, Town05 and Town06, trained in Town03) to validate the generalization performance of the NFRL, SC and Demo modules, as shown in Tables 5. The driving score (DS) served as the primary comprehensive metric, with other indicators providing supplementary reference. The addition of the SC module significantly improves the cross-scenario performance of NFRL (e.g. NFRL vs. NFRL + SC), demonstrating the effectiveness of our SC module design. Further incorporating the Demo learning module on top of NFRL+SC, the experimental results show that NFRL+SC+Demo achieves the highest scores in Town04 (1174.16), Town05 (723.90), and Town06 (2155.40), with substantial improvements over both the baseline NFRL and NFRL+SC configurations. This proves that the Demo module enhances cross-domain generalization through expert knowledge.
As shown in Table 6, we conducted comprehensive comparisons with various mainstream baselines (online RL methods such as L_SAC and Dreamer; offline or imitation learning approaches including BC+Demo and CQL+Demo) across two challenging scenarios (Town03 RND and RBT). The multi-dimensional evaluation metrics clearly demonstrate that: 1) The NFRL framework itself surpasses existing online RL methods; 2) The SC module universally and significantly enhances both safety and overall performance across all methods, including baselines; 3) The NFRL framework effectively utilizes demonstration data, achieving far superior results compared to imitation learning and offline RL baselines; 4) The final NFRL+SC+Demo solution comprehensively outperforms all methods, including enhanced baselines, across nearly all positive metrics (DS, AR, EC, ASD) while maintaining excellent safety performance. These results fully validate the absolute superiority of our proposed method, the effectiveness of each module, and the powerful synergistic effects of their combination.
Table 5: During the evaluation, an ablation study of EGADS’s three modules across scenarios was conducted (Trained in Town03, Evaluated in Town04-Town06, hereinafter referred to as T04-T06)
Method DS $\uparrow$ AR ($f_{1}$) $\uparrow$ EC (%) $\uparrow$ OR (%) $\downarrow$ ASD (m) $\uparrow$
T04 T05 T06 T04 T05 T06 T04 T05 T06 T04 T05 T06 T04 T05 T06
NFRL 326.78 390.54 431.44 1509.90 785.92 947.26 15.81 22.61 29.59 30.88 12.05 16.50 220.18 123.24 143.61
NFRL+SC 649.46 213.17 1234.89 418.22 381.39 1571.85 48.70 38.80 63.42 0.00 0.00 0.00 91.34 50.42 195.36
NFRL+SC+Demo 1174.16 723.90 2155.40 1329.89 894.58 2294.30 57.41 46.85 82.47 3.25 11.51 6.05 159.42 116.96 265.92
4.7 How to improve generalization capabilities ?
Table 6: Evaluation results for different methods in CARLA Town03 (random) and Town03 (roundabout): we denote RND as random and RBT as roundabout. For a fair comparison, all reward functions are in the form of $f_{1}$. Particularly, - indicates that valid data could not be obtained because the episode completion rate for this method is close to 0.
Method DS $\uparrow$ AR ($f_{1}$) $\uparrow$ EC(%) $\uparrow$ OR(%) $\downarrow$ ASD(m) $\uparrow$
RND RBT RND RBT RND RBT RND RBT RND RBT
L_SAC $125.95$ $31.59$ $161.24$ $84.13$ $11.98$ $10.50$ $14.72$ $1.14$ $31.31$ $13.87$
Dreamer 0.03-0.03 0.02-0.02 15.12-15.12 15.12-15.12 $0.01$ $0.02$ - - $0.01$ $0.01$
NFRL 170.03 48.73 424.60 249.17 24.04 10.88 20.93 18.11 72.16 46.27
L_SAC+SC $156.23$ $64.76$ $284.02$ $148.50$ $13.98$ $15.91$ $10.64$ $12.88$ $50.52$ $18.67$
Dreamer+SC $98.12$ $50.02$ $124.74$ $74.98$ $10.42$ $11.20$ $18.08$ $16.35$ $42.85$ $16.90$
NFRL+SC 192.84 101.29 341.56 181.28 38.46 34.66 5.87 4.04 80.21 50.24
BC+Demo 6.30-6.30 1.63-1.63 62.43-62.43 27.92-27.92 $9.22$ $10.31$ $15.49$ $15.57$ $14.78$ $15.34$
CQL+Demo $8.52$ $4.35$ $42.10$ $49.06$ $10.58$ $8.21$ $13.45$ $19.08$ $19.25$ $16.01$
NFRL+Demo $203.26$ $143.03$ $478.04$ $26.15$ $25.71$ $20.66$ $10.50$ $12.82$ $81.32$ $64.80$
NFRL+SC+Demo 485.92 380.17 720.27 653.21 44.25 36.63 7.69 5.48 100.13 84.92
The EGADS system enhances cross-scenario generalization through the co-design of the NFRL framework, SC module, and Demo module. NFRL decouples state representation from policy learning, establishing a transferable foundation for driving policies. As shown in Table 3, in the cross-town evaluation (Town04-Town06), NFRL achieves a significantly higher DS value compared to traditional reinforcement learning methods, demonstrating robust generalization capabilities.
As evidenced in Tables 5 and 6, the SC module effectively mitigates high-risk behaviors through trajectory smoothing, improving overall DS values compared to standalone NFRL and enhancing system robustness. Meanwhile, the Demo module accelerates policy convergence and optimizes exploration via imitation learning. As shown in Tables 5 and 6, the NFRL+SC+Demo configuration demonstrates significant improvements across multiple metrics including DS and AR , confirming that demonstration data effectively reduces inefficient sampling.
The synergy between the SC module and Demo data can be summarized as follows: the SC module establishes safety boundaries to prevent the policy from entering hazardous or suboptimal states, while Demo data alleviates the conservatism of the SC module. EGADS integrates imitation learning (BC loss) and reinforcement learning (NFRL loss), dynamically balancing their weights to enable the agent to leverage expert knowledge while exploring autonomously within safe limits. This balanced mechanism enhances the policy’s generalization capability and environmental adaptability, enabling efficient task execution across diverse scenarios and rapid adaptation to new challenges.
5 Conclusion
In summary, our EGADS framework effectively enhances sample efficiency, safety, and generalization in autonomous driving systems. The inclusion of safety constraints significantly enhances vehicle safety. NFRL, our proposed method, accurately predicts future driving actions, reducing sample complexity. By fine-tuning with a small amount of expert data, NFRL agents learn more general driving principles, which greatly improve generalization and sample complexity reduction, offering valuable insights for autonomous driving system design.
Appendix A Appendix
In the appendix, we provide more details regarding the efficient and generalized end-to-end autonomous driving system with latent deep reinforcement Learning and demonstrations in the paper, including
  • In Subsection A.1, we provide a detailed description of the maps used for training in Gym-Carla.
  • In Subsection A.2, we explain how we collected the expert demonstration dataset in CARLA using a human expert driving with the G29 steering wheel.
  • In Subsection A.3, we explore the impact of up to 12 different data input types on the performance of the NFRL agent.
  • In Subsection A.4, we introduce our reward function with safety constraints.
  • In Subsection A.5, we provide a comprehensive measurement of driving performance metrics.
  • In Subsection A.6, we present the hyperparameter settings for the methods involved in our experiments.
  • In Subsection A.7, we demonstrate the reconstruction of original sensor input images by our NFRL
  • In Subsection A.8, we provide additional results on predictions of future driving actions for NFRL in the imagination space.
A.1 Training CARLA maps